# Impact of electronic correlations on the superconductivity of high-pressure CeH9

### v1

@@ Introduction
[…]
To gain insight into the question of the role of electronic correlation in rare-earth superhydride, we choose CeH9 as a representative example and explore many-body effects using a combination method of DFT and dynamical mean field theory (DMFT), known as DFT+DMFT [25, 26, 27, 28, 29, 30, 31, 32, 33]. Previous studies have shown that the strong correlation nature of cerium $f$-electrons plays a pivotal role in the isostructural phase transition of cerium in its elemental form [34, 35, 36, 37, 38], in cerium intermetallic compounds [39, 40], other rare-earth metals [41, 42], Kondo insulators [43, 44], as well as heavy-fermion systems [45, 46]. We address the notable gap between the experimental results [9, 14, 10, 11] and theoretical forecasts [12, 13, 47] regarding the critical temperature of phonon-mediated superconductivity in CeH9 under high pressure. Theoretical calculations based on DFT underestimate the superconducting critical temperature by nearly 50% [9], prompting speculation about the role of electronic correlations. Our work enhances the description of phonon-mediated superconductivity by incorporating electron-correlation effects for electronic and phononic interactions.
Results and discussion
@@
We study the impact of electronic correlation effects in CeH9 by comparing DFT band-theory results to DMFT many-body calculations. To this end, we first obtain the crystal structure of CeH9 from first principles by optimizing the internal atomic coordinates and the lattice constants of the crystal under 200 GPa using DFT, see Sec. S1 in the Supplementary Information (SI). We find that including the Ce $4f$-electron as valence electrons in the pseudopotential calculations is necessary to obtain lattice parameters in agreement with measurements [48], suggesting that the Ce-$4f$-electrons, although being more closely bound, are not atomic-like. Turning to the electronic structure, we compare in Fig. 1 the DFT band-structure (red lines) to the DMFT spectral function (color map). As in LaH10 [22], the dispersions around the Fermi level are dominated by H $1s$-orbitals that hybridize considerably with the Ce $4f$-orbitals, see Sec. S2 of the SI for additional details. Upon examining the intersections of the (red) bands and the high-intensity spectral features with the Fermi level ($\omega=0$; horizontal white line), we observe that the incorporation of electronic correlation effects results in a Fermi surface that remains largely unaltered in comparison to DFT calculations. In addition, the multiple electron and hole pockets remain qualitatively similar throughout the Brillouin zone. Quantitatively, however, higher-energy states are notably renormalized towards the Fermi level. For example, the electron pocket at the $\bm{\Gamma}$ point becomes 0.4 eV shallower as compared to the DFT result. Notable are the energy shifts of the nondispersive Ce $4f$-orbital bands, originally 1 to 2 eV above the Fermi level, which decrease significantly in DFT + DMFT computations.
@@
Figure 1: |||| Comparison of the electronic bands obtained from DFT and spectral function from DMFT. Momentum-resolved spectral function $A_{\mathbf{k}}(\omega)$ of CeH9 along a high-symmetry path calculated with DMFT (color map) compared to the energy bands calculated by DFT (solid red lines).
Dynamical correlations drive these spectral changes: DFT + U produces a DOS similar to DFT, see Fig. S5 of SI. Meanwhile, the orbital occupation of Ce $4f$ increases slightly from 1.42 in DFT to 1.58 with DFT+DMFT. This charge transfer results in effective hole-doping of the H $1s$-spectral weight. Wang et al. [47] noted that chemical hole doping can shift the electron pocket at $\Gamma$ toward the Fermi level, enhancing the electron-phonon coupling. Remarkably, electronic correlations cause a similar effect in the stoichiometric compound. However, despite broadenings in DFT+DMFT, the excitation spectrum remains coherent and band-like. This can be understood by examining the DFT+DMFT self-energy $\Sigma(\omega)$, from which we derive an averaged local quasiparticle weight $Z=(1-\mathrm{d}\mathfrak{Re}\Sigma(\omega)/\mathrm{d}\omega)^{-1}=0.8$. The imaginary part of the self-energy is negligible near ±$\pm 3$ eV of the Fermi level, allowing a quasiparticle description of the low-energy spectrum. See the Method Section and SI Sec. S3 for more details.
@@
Figure 2: |||| Comparison of density of states (DOS) obtained from DFT and DMFT. a Electronic and b phononic DOS of CeH9 calculated with DFT and DFT+DMFT, respectively. The inset in a highlights the increase in the DOS near the Fermi level.
We now investigate the impact of electronic correlations on the phonon modes. We compute phonon dispersion via the finite difference method [49, 50], at the DFT level using the forces arising from small atomic displacements and at the DFT+DMFT level using forces which are instead defined from the derivatives of the free energy $F=E-TS$, where $T$ is the electronic temperature and $S$ is the impurity entropy [32]. These calculations are enabled by the efficient methodology in Ref. [51]. Figure 2(b) compares the phonon DOS of CeH9 obtained within DFT and DFT + DMFT. Because of the significant difference in the atomic masses, the vibrations of cerium and hydrogen are largely decoupled, resulting in a clear energy separation in the phonon DOS. The low-frequency modes ($\omega<50$ meV) arise predominantly from cerium vibrations, while the high-frequency (optical) modes ($\omega>50$ meV) emerge from dynamical distortions of the hydrogen cages. The highest frequency reaches $240$ meV at the DFT level (red curve), slightly lower than in LaH10 in which the hardest mode reaches approximately 275 meV under similar pressure conditions [52]. The effect of DMFT is to soften hydrogen vibrations.
@@
Finally, we examine the influence of electronic correlations on the superconductivity of CeH9. To contextualize our results, it is worth first clarifying the hierarchy of methods used to evaluate $T_{\mathrm{c}}$: (i) the McMillan-Allen-Dynes (AD) formula ($T^{\mathrm{AD}}_{\mathrm{c}}$[54, 55], which provides a closed-form expression for $T_{\mathrm{c}}$ as follows
@@
where $\omega_{\ln}$ is the logarithmic averaged phonon frequency, $k_{\mathrm{B}}$ denotes the Boltzmann constant, $\lambda$ is the el-ph coupling strength, and $\mu_{\mathrm{c}}$ is the Morel-Anderson pseudopotential accounting for Coulomb repulsion of the system [56]; (ii) the isotropic ($T^{\mathrm{iso}}_{\mathrm{c}}$) and (iii) anisotropic Eliashberg equations ($T^{\mathrm{aniso}}_{\mathrm{c}}$[57, 58], in which $T_{\mathrm{c}}$ is determined implicitly by identifying the highest temperature at which a nonzero superconducting gap $\Delta$ appears. These three methods provide a progressively more refined description of phonon-mediated superconductivity with increasing accuracy and computational cost.
Figure 3: |||| Phonon and electron-phonon properties obtained from DFT and DMFT. Phonon dispersion, Eliashberg spectral function $\alpha^{2}F(\omega)$, and integrated electron-phonon coupling strength $\lambda(\omega)$ as a function of phonon frequency $\omega$ as calculated with a DFT and b DFT+DMFT. The linewidth of the phonon dispersion is proportional to $\lambda(\omega)$ projected onto the phonon modes.
@@
Figure 4: |||| Anisotropic superconducting gap of CeH9. Calculated superconducting gap $\Delta$ on the Fermi surface at 10 K computed with a DFT and b DFT+DMFT. c Anisotropic Migdal-Eliashberg superconducting gap $\Delta$ as a function of temperature. We consider four cases: (top left) DFT electrons with DFT phonons, (bottom left) DMFT electrons with DFT phonons, (top right) DFT electrons with DMFT phonons, and (bottom right) DMFT electrons with DMFT phonons. The solid red line represents the temperature dependence of the superconducting gap expected from the BCS theory [62] in the weak-coupling limit where $\Delta(T)=A\sqrt{1-T/T_{\mathrm{c}}}$, where $A$ is a fitting constant.
Interestingly, with both electrons and phonons described at the DFT level, we find that the el-ph coupling strength for CeH9 is $\lambda=2\int_{0}^{\infty}\mathrm{d}\omega\,\alpha^{2}F(\omega)/{\omega}=0.70$. We then compute the superconducting critical temperatures from the AD formula finding $T^{\mathrm{AD}}_{\mathrm{c}}\simeq 30$ K, and by solving the isotropic and anisotropic Eliashberg equations, leading to $T^{\mathrm{iso}}_{\mathrm{c}}\simeq 37$ K and $T^{\mathrm{aniso}}_{\mathrm{c}}\simeq 47$ K, respectively. As anticipated, the McMillan–Allen–Dynes estimate is the lowest, since that formula is formally valid only for low-$T_{\mathrm{c}}$ superconductors. Our values for $\lambda,$ $T_{\mathrm{c}}^{\mathrm{AD}}$ and $T_{\mathrm{c}}^{\mathrm{iso}}$ agree well with a previous study [9]. The Eliashberg treatment improves the prediction but the anisotropic Eliashberg solution demonstrates that anisotropy significantly increases the predicted $T_{\mathrm{c}}$, bringing it closer to experimental observations.
Figure 3a shows the integrated el-ph coupling strength, $\lambda$($\omega$)=$2\int_{0}^{\omega}\mathrm{d}\omega\,\alpha^{2}F(\omega)/{\omega}$, projected onto the phonon dispersion along with the Eliashberg function $\alpha^{2}F(\omega)$. We find that $\lambda$($\omega$) is distributed in a nonuniform manner among the phonon modes and the phonon Brillouin zone, with approximately 80% of the total el-ph coupling strength arising from the optical phonon modes of hydrogen and the remaining 20% from the phonon modes of cerium. In particular, as highlighted by the linewidth in Fig. 3a, phonon modes with frequency around 125 meV are highly coupled to electrons, leading to a significant peak in the Eliashberg function $\alpha^{2}F(\omega)$. This frequency range is also where we observed a significant softening of the hydrogen mode by electronic correlations.
To go further, we substitute the electron Hamiltonian and phonon dynamical matrices obtained at the DFT level with their counterparts at the DMFT level. These two substitutions will affect the electronic dispersion, the phonon dispersion, and, in principle, the el-ph coupling matrix elements. In this work, we neglect the effect of the latter and use DFT el-ph matrix elements throughout. Current implementations of the Migdal-Eliashberg theory of superconductivity are based on the quasiparticle approximation. Therefore, we construct from our DFT+DMFT calculations an effective quasiparticle band structure for CeH9. The construction of a renormalized band-structure is possible as the real-part of the self-energy can, to good approximation, be linearlized around the Fermi level, while its imaginary part is small. Then, an effective band structure can be found as the solutions $E_{n\mathbf{k}}$ of the linearized quasiparticle equation [63]
$\operatorname{det}\left(E_{n\mathbf{k}}\mathbb{1}-\mathbf{Z}_{\mathbf{k}}\left[\mathbf{H}_{\mathbf{k}}^{\mathrm{\tiny DFT+DMFT}}+\mathfrak{Re}\left\{\bm{\Sigma}_{\mathbf{k}}(0)\right\}-\mu\mathbb{1}\right]\right)=0,$ (2)
which is written in the band basis. Bold quantities are matrices in the band indices $n$, $\mathbb{1}$ is the identity matrix, $\mathbf{Z}_{\mathbf{k}}=\left(1-\left.\mathrm{d}\mathfrak{Re}\bm{\Sigma}_{\mathbf{k}}(\omega)/\mathrm{d}\omega\right|_{\omega=0}\right)^{-1}$ is the quasiparticle weight which approaches unity in a noninteracting system and zero in the limit of extremely strong correlations, $\mathbf{H}_{\mathbf{k}}^{\mathrm{\tiny DFT+DMFT}}=\mathrm{diag}(\ldots,\varepsilon_{n\mathbf{k}},\ldots)$ the Kohn-Sham Hamiltonian at the DFT+DMFT self-consistent charge density, and $\mu$ is the chemical potential.
@@
We validate the quality of the effective band structure obtained using the above quasiparticle approximation in Sec. S4 of the SI. Using this effective band structure and the DMFT phonon spectrum obtained previously, we calculate the el-ph coupling strength and the Eliashberg function at the DMFT level, as shown in Fig. 3b. The total coupling strength in this case is found to increase by nearly 39%, reaching $\lambda\simeq 0.97$. This increment supports a substantial elevation in the critical temperature, which increases from $T_{\mathrm{c}}^{\mathrm{AD}}=30$ K at the DFT level to $T_{\mathrm{c}}^{\mathrm{AD}}=69$ K at the DMFT level. We traced the increase in $T_{\mathrm{c}}$ to changes in specific regions in the Brillouin zone.
@@
Figures 4a-b compare the superconducting gap $\Delta$ on the Fermi surface of CeH9 at 10 K. We find that in the case of DFT the superconducting gap is quite uniform across the Fermi surface, with an average magnitude of approximately 6.3 meV that varies by ±$\pm 2.5$ meV through the Brillouin zone. In contrast, at the DMFT level, the superconducting gap around the corner of the Brillouin zone (i.e. the $K$ point) exhibits a remarkable increase, up to 20 meV, which explains the elevated superconducting critical temperature, from $T_{\mathrm{c}}^{\mathrm{aniso}}=47$ K at the DFT level to $T_{\mathrm{c}}^{\mathrm{aniso}}=96$ K at the DMFT level.
@@
Our results align closely with previous studies for DFT but are underestimated by 50% measurements at comparable pressures. In contrast, our value at the DMFT level, $T_{\mathrm{c}}^{\mathrm{aniso}}=96$ K at 200 GPa, is in excellent agreement with the experimental value, $T_{\mathrm{c}}=95$ K at the closest pressure of 150 GPa. Dissecting the influence of electronic correlations on the different ingredients entering the calculation of $T_{\mathrm{c}}$, we find that the DMFT renormalization of the electronic structure is the main driving force behind the increased critical temperature. Indeed, keeping the phonons at the DFT-level, the effect of DMFT on electrons almost doubles the $T_{\mathrm{c}}$. Indeed, as shown in Fig 2a, the increased DOS at the Fermi level, $D(E_{\mathrm{F}})$, facilitates the formation of more Cooper pairs, leading to a higher superconducting critical temperature [62] for which $T_{\mathrm{c}}\propto\exp{(1/VD(E_{\mathrm{F}}))}$ with $V$ the average el-ph coupling constant. Finally, we also discover that exclusively renormalizing the electronic structure with DMFT still leads to a substantial underestimation of the superconducting critical temperature. It is necessary to also include the phonon-softening effect owing to electronic correlations to quantitatively reconcile the theoretical findings with the experimental data.
Conclusion

### v2

@@ Introduction
Among the various candidates, rare-earth metal superhydrides have shown great promise, especially with the discovery of LaH10 [5, 6, 7, 8]. It exhibits a superconducting critical temperature of around 250 K at 170 GPa, marking the highest measured superconducting critical temperature to date. Furthermore, cerium superhydrides (CeHn with $n=9$ or $10$) have also attracted attention: although their maximum superconducting critical temperatures do not exceed that of LaH10, they reach a relatively high superconducting critical temperature (around $100$ K) at pressures on the order of 100 GPa or less [9, 10, 11]. This lower‐pressure regime not only simplifies high‐pressure synthesis and characterization but also further brings the goal of ambient‐pressure superconductivity closer to fruition. In general, numerous rare-earth superhydrides have been predicted and synthesized in the past decade [12, 13, 14, 15, 16, 17, 5, 9, 18, 19, 20], but a pivotal yet unresolved question persists. What role do electronic correlations, induced by the $f$-electrons of the rare-earth elements, play in these high-pressure compounds? In LaH10, it has been observed that external pressure destabilizes the 6ss- and 5dd-orbitals of La, making the $4f$-orbitals energetically more favorable [21]. As a result, the electronic states of LaH10 near the Fermi level are composed of La $4f$- and H $1s$-orbitals [22]. We note that these states involving La $4f$-orbitals are strongly localized and generally cannot be well described by density functional theory (DFT) [23, 24].
To gain insight into the question of the role of electronic correlation in rare-earth superhydride, we choose CeH9 as a representative example and explore many-body effects using a combination method of DFT and dynamical mean field theory (DMFT), known as DFT+DMFT [25, 26, 27, 28, 29, 30, 31, 32, 33]. Previous studies have shown that the strong correlation nature of cerium $f$-electrons plays a central role in the isostructural phase transition of cerium in its elemental form [34, 35, 36, 37, 38], in cerium intermetallic compounds [39, 40], other rare-earth metals [41, 42], Kondo insulators [43, 44], as well as heavy-fermion systems [45, 46]. We address the notable gap between the experimental results [9, 14, 10, 11] and theoretical forecasts [12, 13, 47] regarding the critical temperature of phonon-mediated superconductivity in CeH9 under high pressure. Theoretical calculations based on DFT underestimate the superconducting critical temperature by nearly 50% [9], prompting speculation about the role of electronic correlations. Our work enhances the description of phonon-mediated superconductivity by incorporating electron-correlation effects for electronic and phononic interactions.
Results
@@
We study the impact of electronic correlation effects in CeH9 by comparing DFT results to DMFT many-body calculations with the nominal double-counting scheme [37]. To this end, we first obtain the crystal structure of CeH9 from first principles by optimizing the internal atomic coordinates and the lattice constants of the crystal under 200 GPa using DFT, see Sec. S1 in the Supplementary Information (SI). Treating the Ce $4f$-electron as valence electrons in the pseudopotential calculations is necessary to obtain lattice parameters in agreement with measurements [48], suggesting that the Ce-$4f$-electrons, although being more closely bound, are not atomic-like. Turning to the electronic structure, we compare in Fig. 1 the DFT band-structure (red lines) to the DMFT spectral function (color map). As in LaH10 [22], the dispersions around the Fermi level are dominated by H $1s$-orbitals that hybridize considerably with the Ce $4f$-orbitals, see Sec. S2 of the SI for additional details. Upon examining the intersections of the red bands and the high-intensity spectral features with the Fermi level (horizontal white line), we observe that the incorporation of electronic correlation effects results in a Fermi surface that remains largely unaltered in comparison to DFT calculations. In addition, the multiple electron and hole pockets remain qualitatively similar throughout the Brillouin zone. Quantitatively, however, higher-energy states are notably renormalized towards the Fermi level. For example, the electron pocket at the $\bm{\Gamma}$ point becomes 0.4 eV shallower as compared to the DFT result. Notable are the energy shifts of the nondispersive Ce $4f$-orbital bands, originally 1 to 2 eV above the Fermi level, which decrease significantly in DFT + DMFT computations.
Figure 1: $|$ Comparison of the electronic bands obtained from DFT and spectral function from DMFT. Momentum-resolved spectral function $A_{\mathbf{k}}(\varepsilon)$ of CeH9 along a high-symmetry path calculated with DMFT (color map) compared to the energy bands calculated by DFT (solid red lines).
@@
We also find that DFT + U, using the fully-localized-limit double counting scheme, produces a DOS similar to DFT, see Fig. S5 of SI, suggesting that dynamical correlations drive the observed spectral changes. However, differences in the double-counting scheme between DFT+U and DFT+DMFT may also play a role. Meanwhile, the orbital occupation of Ce $4f$ increases slightly from 1.42 in DFT to 1.58 with DFT+DMFT. This charge transfer results in effective hole-doping of the H $1s$-spectral weight. Wang et al. [47] noted that chemical hole doping can shift the electron pocket at $\bm{\Gamma}$ toward the Fermi level, enhancing the electron-phonon coupling. Remarkably, electronic correlations cause a similar effect in the stoichiometric compound. However, despite broadenings in DFT+DMFT, the excitation spectrum remains coherent and band-like. This can be understood by examining the local DFT+DMFT self-energy $\Sigma_{L}(\varepsilon)$, from which we derive an averaged local quasiparticle weight $Z=\frac{1}{N_{L}}\sum_{L}\big[1-\frac{\partial\Re\Sigma_{L}(\varepsilon)}{\partial\varepsilon}\big]_{\varepsilon=\varepsilon^{\rm F}}=0.8$, where $L$ is the $4f$ manifold with $N_{L}=7$. The imaginary part of the self-energy is negligible around $\pm 3$ eV of the Fermi level, allowing a quasiparticle description of the low-energy spectrum. See the Method Section and SI Sec. S3 for more details.
@@
Figure 2: $|$ Comparison of density of states (DOS) per unit cell obtained from DFT and DMFT. a Electronic and b phononic DOS of CeH9 calculated with DFT and DFT+DMFT, respectively. The inset in a highlights the increase in the DOS near the Fermi level.
We now investigate the impact of electronic correlations on the phonon modes. We compute phonon dispersion via the finite difference method [49, 50], at the DFT level using the forces arising from small atomic displacements and at the DFT+DMFT level using forces which are instead defined from the derivatives of the free energy F=ETSF=E-TS, where $T$ is the electronic temperature and $S$ is the impurity entropy [32]. These calculations are enabled by the efficient methodology from Ref. [51]. Figure 2b compares the phonon DOS of CeH9 obtained within DFT and DFT + DMFT. Because of the significant difference in the atomic masses, the vibrations of cerium and hydrogen are largely decoupled, resulting in a clear energy separation in the phonon DOS. The low-frequency modes ($\omega<50$ meV) arise predominantly from cerium vibrations, while the high-frequency modes ($\omega>50$ meV) emerge from dynamical distortions of the hydrogen cages. The highest frequency reaches $240$ meV at the DFT level (red curve), slightly lower than in LaH10 in which the hardest mode reaches approximately 275 meV under similar pressure conditions [52]. The effect of DMFT is to soften hydrogen vibrations.
@@
Finally, we examine the influence of electronic correlations on the superconductivity of CeH9. To contextualize our results, it is worth first clarifying the hierarchy of methods used to evaluate the superconducting transition temperature ($T_{\mathrm{c}}$): (i) the McMillan-Allen-Dynes (AD) formula ($T^{\mathrm{AD}}_{\mathrm{c}}$) [54, 55], which provides a closed-form expression for $T_{\mathrm{c}}$ as follows
Figure 3: $|$ Phonon and electron-phonon properties obtained from DFT and DMFT. Phonon dispersion, Eliashberg spectral function $\alpha^{2}F(\omega)$, and integrated electron-phonon coupling strength $\lambda(\omega)$ as a function of phonon frequency $\omega$ as calculated with a DFT and b DFT+DMFT. The linewidth of the phonon dispersion is proportional to $\lambda(\omega)$ projected onto the phonon modes.
@@
Figure 4: $|$ Anisotropic superconducting gap of CeH9. Calculated superconducting gap $\Delta$ on the Fermi surface at 10 K computed with a DFT and b DFT+DMFT. c Anisotropic Migdal-Eliashberg superconducting gap $\Delta$ as a function of temperature. We consider four cases: (top left) DFT electrons with DFT phonons, (bottom left) DMFT electrons with DFT phonons, (top right) DFT electrons with DMFT phonons, and (bottom right) DMFT electrons with DMFT phonons. The solid red line represents the temperature dependence of the superconducting gap expected from the BCS theory [56] in the weak-coupling limit where $\Delta(T)=A\sqrt{1-T/T_{\mathrm{c}}}$, where $A$ is a fitting constant.
@@
where $\omega_{\ln}$ is the logarithmic averaged phonon frequency, $k_{\mathrm{B}}$ denotes the Boltzmann constant, $\lambda$ is the electron-phonon (el-ph) coupling strength, and $\mu_{\mathrm{c}}$ is the Morel-Anderson pseudopotential accounting for Coulomb repulsion of the system [58]; (ii) the isotropic ($T^{\mathrm{iso}}_{\mathrm{c}}$) and (iii) anisotropic Eliashberg equations ($T^{\mathrm{aniso}}_{\mathrm{c}}$) [59, 60], in which $T_{\mathrm{c}}$ is determined by identifying the highest temperature at which a nonzero superconducting gap $\Delta$ appears. These three methods provide a progressively more refined description of phonon-mediated superconductivity with increasing accuracy and computational cost.
Interestingly, with both electrons and phonons described at the DFT level, we find that the el-ph coupling strength for CeH9 is $\lambda=2\int_{0}^{\infty}\mathrm{d}\omega\,\alpha^{2}F(\omega)/{\omega}=0.70$. We then compute the superconducting critical temperatures from the AD formula finding $T^{\mathrm{AD}}_{\mathrm{c}}\simeq 30$ K, and by solving the isotropic and anisotropic Eliashberg equations, leading to $T^{\mathrm{iso}}_{\mathrm{c}}\simeq 37$ K and $T^{\mathrm{aniso}}_{\mathrm{c}}\simeq 47$ K, respectively. As anticipated, the McMillan–Allen–Dynes estimate is the lowest, since that formula is formally valid only for low-$T_{\mathrm{c}}$ superconductors. Our values for $\lambda,$ $T_{\mathrm{c}}^{\mathrm{AD}}$ and $T_{\mathrm{c}}^{\mathrm{iso}}$ agree well with a previous study [9]. The Eliashberg treatment improves the prediction. Importantly, the anisotropic Eliashberg solution demonstrates that anisotropy significantly increases the predicted $T_{\mathrm{c}}$, bringing it closer to experimental observations.
[…]
To go further, we substitute the electron Hamiltonian and phonon dynamical matrices obtained at the DFT level with their counterparts at the DMFT level. These two substitutions affect the electronic dispersion, the phonon dispersion, and the el-ph coupling matrix elements. We update the latter via the DMFT phonon eigendisplacement and frequencies but we neglect the effect of DMFT on the perturbed electronic potential. Current implementations of the Migdal-Eliashberg theory of superconductivity are based on the quasiparticle approximation. Therefore, we construct from our DFT+DMFT calculations an effective quasiparticle band structure for CeH9. The construction of a renormalized band-structure is possible as the real-part of the self-energy can, to good approximation, be linearlized around the Fermi level where its imaginary part is small. Then, an effective band structure can be found by upfolding the local DMFT self-energy into the Kohn-Sham basis $\Sigma_{nn^{\prime}\mathbf{k}}(\varepsilon)=\sum_{L}P_{nn^{\prime}L\mathbf{k}}\Sigma_{L}(\varepsilon)$ where $P_{nn^{\prime}L\mathbf{k}}$ is the inverse of the projector that defines the local $4f$ basis [31]. In the DFT+DMFT, the charge density is obtained self-consistently, leading to an updated diagonal Kohn-Sham Hamiltonian $H^{\rm DFT+DMFT}_{nn\mathbf{k}}$. We obtain the renormalized quasiparticle energies $E_{n\mathbf{k}}$ by solving the following linearized equation [63]:
$\operatorname{det}\left(E_{n\mathbf{k}}\delta_{nn^{\prime}}-\sum_{m}Z_{nm\mathbf{k}}\left[H_{mn^{\prime}\mathbf{k}}^{\mathrm{\tiny DFT+DMFT}}+\Re\Sigma_{mn^{\prime}\mathbf{k}}(\varepsilon^{\rm F})\right]\right)=0,$ (2)
where
$Z_{nn^{\prime}\mathbf{k}}=\bigg[1-\frac{\partial\Re\Sigma_{nn^{\prime}\mathbf{k}}(\varepsilon)}{\partial\varepsilon}\Big|_{\varepsilon=\varepsilon^{\rm F}}\bigg]^{-1}$ (3)
is the quasiparticle weight which approaches unity in a noninteracting system and zero in the limit of extremely strong correlations.
@@
We validate the quality of the effective band structure obtained using the above quasiparticle approximation in Sec. S4 of the SI. Using this effective band structure and the DMFT phonon spectrum obtained previously, we calculate the el-ph coupling strength and the Eliashberg function at the DMFT level, as shown in Fig. 3b. The total coupling strength in this case is found to increase by nearly 39%, reaching $\lambda\simeq 0.97$. This increase supports a substantial elevation in the critical temperature, which increases from $T_{\mathrm{c}}^{\mathrm{AD}}=30$ K at the DFT level to $T_{\mathrm{c}}^{\mathrm{AD}}=69$ K at the DMFT level. We traced the increase in $T_{\mathrm{c}}$ to changes in specific regions in the Brillouin zone.
@@
Figures 4a-b compare the superconducting gap $\Delta$ on the Fermi surface of CeH9 at 10 K. We find that in the case of DFT the superconducting gap is quite uniform across the Fermi surface, with an average magnitude of approximately 6.3 meV that varies by $\pm 2.5$ meV through the Brillouin zone. In contrast, at the DMFT level, the superconducting gap around the corner of the Brillouin zone (i.e. the $K$ point) exhibits a remarkable increase, up to 20 meV, which explains the elevated superconducting critical temperature, from $T_{\mathrm{c}}^{\mathrm{aniso}}=47$ K at the DFT level to $T_{\mathrm{c}}^{\mathrm{aniso}}=96$ K at the DMFT level.
@@
Our results align closely with previous DFT studies but overestimate experimental values at comparable pressures by 50%. In contrast, our value at the DMFT level, $T_{\mathrm{c}}^{\mathrm{aniso}}=96$ K at 200 GPa, is in excellent agreement with the experimental value, $T_{\mathrm{c}}=95$ K at the closest pressure of 150 GPa. Dissecting the influence of electronic correlations on the different ingredients entering the calculation of $T_{\mathrm{c}}$, we find that the DMFT renormalization of the electronic structure is the main driving force behind the increased critical temperature. Indeed, keeping the phonons at the DFT-level, the effect of DMFT on electrons almost doubles the $T_{\mathrm{c}}$. Indeed, as shown in Fig 2a, the increased DOS at the Fermi level, $D(\varepsilon^{\mathrm{F}})$, facilitates the formation of more Cooper pairs, leading to a higher superconducting critical temperature for which $T_{\mathrm{c}}\propto\exp{(-\frac{1}{VD(\varepsilon^{\mathrm{F}})})}$ where $V$ is the pairing interaction [56]. Finally, we also discover that exclusively renormalizing the electronic structure with DMFT still leads to a substantial underestimation of the superconducting critical temperature. It is necessary to also include the phonon-softening effect due to electronic correlations to quantitatively reconcile the theoretical findings with experimental data.
Discussion